Sol Summary – January 13th


 Crew 219 Sol Summary Report 13-01-2020

Sol: 8

Summary Title: Oh, Mondays...

Author's name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: The crew was woken abruptly around 0600 to a
generator failure and low power, and immediately responded according
to guidance from Outpost. We then boiled water on the stove for
coffee, and got started with the days work. Some worked on science
while others transplanted plants in the GreenHab. H. Blackburn made a
lunch of chili and rice before EVA #13 pushed out towards Beige Moon,
and those left behind continued their tasks. EVA #13 collected their
samples, but suffered an H&S-related incident (see HSO report) forcing
their return by 1508. K. Crisman resolved the situation swiftly,
decisively, and admirably. With the whole crew back at nominal status,
A. Elnajdi and R. Raphael prepared an incredible dinner for the crew
while everyone finished and proofread reports for this evening’s Comms

Look Ahead Plan: One EVA, continued sample analysis, and maintenance
as necessary

Anomalies in work: Generator Failure

Weather: Early morning fog, cold and sunny all day

Crew Physical Status: Nominal at present (See HSO Report)


EVA #13 – Geologic Sample Collection around Beige Moon (0516000/4254300)

Reports to be filed:

Sol Summary

EVA Report #13

EVA Request #14

HSO Report

Operations Report

GreenHab Report

Research Report

Journalist's Report

Photos (Including Photo of the Day)

Support Requested:

-Quart-sized ziploc bags

-Medical tape

-Refer to Operations Report for maintenance-related request 
Copyright © The Mars Society. All rights reserved. | Main Site